OpenStudy (anonymous):

An isosceles triangle in which angles B and C have the same measure is shown. Find the measure of each angle whose degree measure us represented with variables. Angle A: x+8y-1 Angle B: 3y+4 Angle C:7x+5

OpenStudy (anonymous):

3y + 4 = 7x + 5 and x + 8y - 1 + 3y + 4 + 7x + 5 + 180

OpenStudy (anonymous):

x + 8y - 1 + 3y + 4 + 7x + 5 = 180

OpenStudy (anonymous):

y = -4/11 (-43+2 x)

OpenStudy (anonymous):

x = 5

OpenStudy (anonymous):

3y + 4 = 35 + 5 3y = 36 y = 12

OpenStudy (anonymous):

that is right.

OpenStudy (anonymous):

angle A is 100 degrees angle B is 40 degrees angle C is 40 degrees

OpenStudy (anonymous):

thanks, now let me study the steps

OpenStudy (anonymous):

genius

OpenStudy (anonymous):

moneybird come back

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

not

OpenStudy (anonymous):

It's given that B = C, right?

OpenStudy (anonymous):

x + 8y - 1 + 3y + 4 + 7x + 5 = 180 show work please

OpenStudy (anonymous):

the sum of all angles in a triangle is equal to 180

OpenStudy (anonymous):

true

OpenStudy (anonymous):

collect like terms 8x + 11y + 8 = 180 Isolate one variable first

OpenStudy (anonymous):

then substitute that into the equation 3y + 4 = 7x + 5